# VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

### v1

@@ 5.5.1 Set-Transformer Genetic Variant Importance
4217696_TGGAG>>TTGAA 2.594 papA4 (novel)
1262193_C>>T 2.589 pks13 (novel)
The analysis demonstrated the framework’s superior capability to identify both canonical and non-canonical resistance mechanisms. As expected, rpoB coding region variants ranked most highly, with the top-scoring variant 761140_A>T (Score 6.896) confirming the model’s alignment with the shared target mechanism between RFB and RIF (williams1998contribution; van2013rifampin; wang2023antimicrobial). Multiple additional rpoB variants achieved high scores, including 761139_C>A (Score 3.955), 761140_A>G (Score 3.321), 761110_A>G (Score 3.316), and 761110_A>T (Score 2.463), all clustering within the Rifampicin Resistance-Determining Region (RRDR).
Furthermore, the framework successfully captured a Multi-Drug Resistance (MDR) genomic signature from a single drug prediction task. The detection of relevant variants in embB (4247431_G>T) and gyrA (7581_G>A) confirms the learned importance of co-selection and cross-resistance patterns (ghosh2020survey).
The discovery potential of the framework is underscored by the high ranking of numerous novel variants. High scores for variants in genes such as mmpL5 (2944085_C>T, Score 4.947), lpqS (893702_GAC...CCG, Score 4.213), ppsD (3640407_G>A, Score 3.532) reinforce the strong suggestion that these uncharacterized loci represent crucial epistatic interaction points or rare, high-effect genes that contribute significantly to the RFB phenotype. These variants are now prime candidates for functional validation to uncover new biological resistance pathways.
5.5.2 Genetic Variant Interaction Network Analysis
Analysis of the aggregated Set Attention Block (SAB) weights culminates in the Multi-locus Attribution-Weighted Interaction (MAWI) Map (Figures 11 and 12), which provides a critical, quantifiable decomposition of complex, multi-locus epistatic effects underlying Rifampicin (RIF) and Rifabutin (RFB) resistance. This approach moves beyond single-variant association models to capture the necessary synergistic coordination between mutations. Specifically, the heatmaps visualize the highest magnitude of attention-weights found by the SAB transformer, thereby mapping the strength and direction of these interactions and providing mechanistic insights into the polygenic architecture of drug resistance.
Figure 11: Variant interaction heatmaps showing attention-based relationships between top hub variants for RIF resistance. Red indicates strong interactions; yellow indicates weak interactions. The patterns reveal how variants work together in resistance mechanisms
The variant interaction heatmap (Figure 11) for RIF (Rifampicin) resistance reveals a highly interconnected, yet structured, network of epistatic relationships among the top hub variants, with a high overall signal density indicating numerous strong-to-moderate synergistic interactions. This high-interaction density suggests that robust resistance relies on compensatory or cooperative pathways, confirming that the model has effectively learned the structural dependencies of a highly interconnected resistance mechanism. The pattern is characterized by specific "hub" variants, such as 3829770_T>>C and 1307598_C>>G, which act as central nodes, exhibiting dark red cells that indicate strong attention to multiple other variants in the network. This hub-driven structure supports the idea that the mechanism relies on a few key "linchpin" mutations that coordinate with a set of secondary, supportive mutations, and this strong cross-attention quantitatively supports the hypothesis of epistatic fitness compensation. Furthermore, the scattered distribution of high-value cells across the map, rather than being localized to a single cluster, suggests a diffuse polygenic mechanism where many different combinations of mutations contribute significantly to the resistance phenotype.
Figure 12: Variant interaction heatmaps showing attention-based relationships between top hub variants for RFB resistance. Red indicates strong interactions; blue indicates weak interactions. The patterns reveal how variants work together in resistance mechanisms
The second heatmap in Figure 12 (blue-to-red scale, max interaction $\approx 0.40$) depicts a contrasting, sparse network of high-magnitude, localized epistatic effects for RFB resistance. In contrast to Figure 11, this visualization shows a predominantly low-interaction background (dark blue), highlighting a few critical, high-leverage interactions that dominate the attention landscape involving variants like 1070702_T>>C. The high attention-weights (dark red) pattern suggests that these specific variant interactions act as a primary mechanism for resistance, where the effect of one variant is highly dependent on the presence of the other. The sparsity indicates that, while many variants are present, only a select few combinations are mechanistically determinative, allowing the VAMP-Net model to pinpoint the most non-additive synergistic effects.
The contrasting patterns between the two drugs demonstrate that resistance involves coordinated resistance modules whose composition varies depending on the specific drug, providing a key insight for designing combination therapies. The MAWI provides an interpretable, quantitative visualization of complex genetic epistasis, confirming known biological interactions while highlighting high-impact, potentially novel interaction hubs.
5.5.3 VCF Feature Channel Analysis: Quantifying the Influence of Variant Quality Metrics
We quantified the contribution of the eight VCF features via systematic ablation, validating the unique multi-channel design of the Quality-Aware 1D-CNN pathway (Figure 13). This analysis highlights the technical robustness required for reliable genotype-phenotype association.
The analysis confirms the model does not rely on simple variant presence (GT field) but critically leverages the contextual quality and support metrics:
@@ 5.5.3 VCF Feature Channel Analysis: Quantifying the Influence of Variant Quality Metrics
  • Dominant Metric (FRS): The Fraction of Supporting Reads (FRS) is the single most influential feature for both the RIF and RFB models. This strong reliance confirms that the 1D-CNN pathway effectively learns to filter out false positives or low-support variants, as high FRS is used as a primary filter for calling true clonal mutations in MTB sequencing (cryptic2022data).
@@
  • Contextual Confidence: The GT_CONF_PERCENTILE is the second most important feature. By utilizing the relative confidence rank rather than the raw GT_CONF (which shows minimal contribution), the model demonstrates its capacity to leverage comparative, sample-specific context for prediction.
Figure 13: VCF feature importance comparison for RIF and RFB resistance models. Fraction of Supporting Reads (FRS) and Genotype Confidence Percentile (GT_CONF_PERCENTILE) demonstrate substantially higher importance for both resistance phenotypes, with the RFB model showing greater dependence on these quality metrics
The near-zero composite scores for features like Read Depth (DP), Genotype (GT), and Raw GT_CONF are equally informative. The low importance of DP suggests its information is effectively captured by the derived, normalized metrics (FRS and GT_CONF_PERCENTILE), rendering the raw depth redundant. This indicates that the presence of a variant or its raw confidence is insufficient; the model requires the granular, multi-dimensional context of the variant’s support and relative quality to make high-confidence resistance predictions.
@@
This ablation study confirms that the 1D-CNN pathway successfully extracts biologically and technically meaningful information from the VCF features, enabling a robust distinction between true resistance-conferring variants and lower-quality calls, especially critical for subtle phenotypes like RFB resistance.
5.5.4 Dual-Path Fusion Model Interpretability
This final analysis validates the Multi-Path Fusion strategy as the core functional advancement of VAMP-Net, establishing its ability to combine Pathogenicity (via the Set-Transformer) and Reliability (via the 1D-CNN).
The analysis revealed two novel findings critical to the framework’s superior performance and generalizability:
  • Learned Adaptive Feature Regularization: The 1D-CNN pathway functions as a learned signal-to-noise regulator, performing non-linear, context-aware feature selection (prioritizing FRS and GT_CONF_PERCENTILE over redundant features like DP). The most significant evidence of this is the 8 to 14-fold increase in feature reliance observed in the RFB model, which demonstrates that the fusion mechanism dynamically compensates for the weaker, more diverse genetic signal of that phenotype by demanding greater technical certainty. Path-2 thus acts as an adaptive attention mechanism that modulates prediction confidence based on data quality.
  • Epistatic Network Auditing: The SAB network analysis confirmed that the final prediction is governed by an epistatic network audit, where the model learns that a high-importance variant from Path-1 will have its final impact dynamically down-weighted if Path-2 reports low sequencing support. This integrated approach provides an auditable, high-confidence score that addresses the critical challenge of separating biological signal from computational support in clinical genomic data.
In conclusion, integrating VCF features in this manner addresses a major challenge in genomic deep learning: the inherent separation between the biological signal (variant identity) and the computational support (variant quality).

### v2

@@ 4.5.1 Mechanistic Interpretation and Variant Attribution
EMB 740348_TCGCGCGAC>>T 0.756 lipG (novel)
764509_G>>C 0.638 rpoC (novel)
740383_G>>T 0.559 lipG (novel)
2435103_T>>C 0.536 idsA2 (novel)
4269298_A>>G 0.180 ubiA (known)
4247429_A>>C 0.062 embB (known)
4247429_A>>T 0.055 embB (known)
4248003_A>>G 0.046 embB (known)
2603465_C>>G 0.502 intergenic (regulatory candidate)
2113033_G>>C 0.363 intergenic (regulatory candidate)
INH 2760136_TCC…ATA>>TTC…ATG 8.195 mmuM (novel)
4366246_GCC…CG>>GGC…GCC 0.763 eccA2 (novel)
575699_G>>A 0.085 mshA (known)
413294_ACCC…>>A 0.097 iniC (known)
1503392_ACATG…GGTTG>>A 0.527 intergenic (regulatory candidate)
1113698_CACG…GATGA>>C 0.502 intergenic (regulatory candidate)
VAMP-Net assigned primary predictive weight to canonical mutations within the RRDR of the $rpoB$ gene, precisely aligning with WHO-validated markers for Rifampicin (RIF) and Rifabutin (RFB), where the top-ranked RFB variant 761140_A>>T achieved a dominant score of 6.896. Notably, the model extracted complex Multi-Drug Resistance (MDR) signatures from single-drug tasks; for RIF, the highest importance score (0.446) was attributed to the Isoniazid-associated $katG$ variant, while for Ethambutol (EMB), the model correctly prioritized the $rpoC$ variant (0.638). This confirms that VAMP-Net leverages co-selection patterns and compensatory mechanisms inherent in MDR-TB lineages to enhance its predictive confidence.A critical outcome of the attribution analysis is the identification of high-scoring variants in uncharacterized regions, most significantly the mmuM variant for Isoniazid (INH), which attained the study’s highest overall importance score of 8.195. Other key discoveries include $ponA1$ (0.308) and $pknB$ (0.273) for RIF, $mmpL5$ (4.947) for RFB, and the novel $lipG$ (0.756) and $idsA2$ (0.536) loci for EMB. The consistent prioritization of these novel loci across all four drugs suggests that VAMP-Net’s attention mechanism successfully isolates rare, high-effect mutations and epistatic modifiers that are often missed by standard frequency-based methods.
4.5.2 Functional Enrichment
To determine whether the high-attribution features identified by VAMP-Net reflect genuine biological mechanisms, we cross-referenced the top-ranked variants against established proteomic and molecular literature. As detailed in Table 16, our framework successfully identified a hierarchical resistance signature that spans from canonical target modification to global cell-envelope remodeling. VAMP-Net correctly prioritized rpoB variants, specifically distinguishing the unique structural binding affinities of Rifabutin (RFB) over Rifampicin (RIF) through the selection of variants like 761140_A>>T (IG: 6.896) [21]. Beyond canonical targets, the framework isolated broad-spectrum efflux mechanisms, notably the mmpL5 system (IG: 4.947), which likely modulates RFB susceptibility through active drug extrusion [18]. Additionally, the high ranking of the persistence-related lpqS gene (IG: 4.213) suggests the model captures the evolutionary trajectory of resistance where enhanced macrophage survival precedes the accumulation of primary resistance markers [37] Significantly, the model identified an integrated metabolic program for cell-wall reinforcement. High-attribution scores for ppsD, fadD26, and papA4, members of a co-transcribed hub, reflect the biosynthesis of the Phthiocerol Dimycocerosates (PDIM) waxy coat [2, 43]. The prioritization of these loci, known to be upregulated in rpoB mutants, demonstrates that VAMP-Net has internalized the compensatory permeability barriers employed by M. tuberculosis under drug pressure. These results confirm that the framework reconstructs a biologically grounded, multi-layered defense strategy rather than relying on arbitrary statistical correlations.
Table 16: Predictive attribution and biological functional validation of representative Rifabutin (RFB) resistance markers identified by VAMP-Net.
Gene Variant IG Score Category Functional Mechanism Reference
rpoB 761140_A>>T 6.896 Primary Canonical RNA Polymerase target; different structural [21]
761139_C>>A 3.955 Resistance affinities for RFB vs. RIF yield distinct mutation profiles.
mmpL5 2944085_C>>T 4.947 Efflux Pump Broad-spectrum efflux (Bedaquiline/Clofazimine); [18]
likely contributes to non-specific RFB resistance.
lpqS 893702_GAC… 4.213 Persistence Resists copper toxicity in macrophages; enhances [37]
persistence and subsequent mutation accumulation.
ppsD 3640407_G>>A 3.532 Cell Wall Essential for PDIM (waxy coat) biosynthesis; directly [2]
(PDIM) upregulated in rpoB mutants to enhance permeability barrier.
fadD26 3545402_A>>G 3.527 Cell Wall Initiates PDIM biosynthesis; co-transcribed with ppsD [43]
(PDIM) to reinforce the outer waxy envelope.
pks13 1262023_C>>T 2.885 Cell Wall Final catalyst for mycolic acid synthesis; primary [15]
1262193_C>>T 2.589 (Mycolic) barrier layer essential for natural drug resistance.
papA4 4217696_TGG.. 2.594 Cell Wall Anchors PDIM to the surface; loss of function leads [19]
(PDIM) to envelope mislocalization and drug vulnerability.
To validate the biological coherence of the identified novel resistance loci, we performed functional enrichment analysis using STRING-db 15. The high-attribution genes (ponA1, pknB, mmpL5, lpqS, ppsD, fadD26, pks13, papA2) formed a statistically significant interaction network ($p=0.00239$), confirming they constitute a biologically non-random module. Gene Ontology (GO) analysis identified six enriched biological processes converging on cell wall remodeling as a unified adaptive strategy: Significant enrichment in "Cell wall organization or biogenesis" ($FDR=7.16\times 10^{-5}$) and "Peptidoglycan-based cell wall biogenesis" ($FDR=0.00056$) directly validates the RIF-associated loci ponA1 and pknB; enrichment in "DIM/DIP cell wall layer assembly" ($FDR=0.0076$) and "Phthiocerol biosynthetic process" ($FDR=0.0257$) confirms that the model autonomously identified the precise PDIM/phthiocerol lipid pathways. High enrichment strength for the "Phthiocerol biosynthetic process" (strength = 2.22) biochemically confirms the functional assembly of ppsD, fadD26, and pks13.These results demonstrate that VAMP-Net‘s discovery reflects the coordinated upregulation of PDIM biosynthesis and cell wall reinforcement as an adaptive epistatic response to rpoB-mediated resistance.
Figure 15: Functional protein-protein interaction network and GO enrichment of novel loci identified for RIF/RFB resistance. The interaction map (top) highlights two significant metabolic modules discovered via VAMP-Net‘s attention weights: the ppsD-pks13-fadD26 cluster and the ponA1-pknB axis. The Gene Ontology analysis (bottom) confirms these loci are statistically enriched for cell wall organization ($p=7.16\times 10^{-5}$) and PDIM biosynthesis ($p=0.0257$). These results demonstrate VAMP-Net’s ability to recover biologically cohesive compensatory mechanisms alongside canonical resistance markers.
Functional enrichment analysis of EMB-associated loci yielded the study‘s strongest validation result 16, centered on Arabinosyltransferase activity (GO:0052636). This term achieved an exceptional enrichment signal of 4.0 ($FDR=3.44\times 10^{-7}$), capturing 4 out of 4 known arabinosyltransferase-encoding genes in the M. tuberculosis genome.This constitutes a near-perfect biological validation, as ethambutol (EMB) specifically targets the arabinosyltransferase enzymes, encoded by embA, embB, and embC—responsible for cell wall synthesis. VAMP-Net autonomously recovered this entire enzyme family along with the ubiA locus, a known resistance gene providing the necessary substrate. Furthermore, "Cell wall organization" (GO:0071555) was significantly enriched ($FDR=0.0053$), consistent with EMB’s role in disrupting cell wall assembly. These findings confirm that the model learns integrated resistance pathways, comprising primary targets, substrate supply, and potential epistatic modifiers, rather than isolated variant associations.
Figure 16: Functional interaction network and GO enrichment for Ethambutol (EMB) resistance loci. The interaction map (top) illustrates a tightly coupled module comprising embA, embB, embC, and ubiA, which are primary drivers of EMB resistance. Gene Ontology analysis (bottom) confirms high statistical enrichment for arabinosyltransferase activity ($p=3.44\times 10^{-7}$) and cell wall organization ($p=0.0053$). These results validate VAMP-Net’s ability to precisely identify the specific metabolic pathways targeted by the drug.
4.5.3 Epistatic Interaction Network Analysis
Analysis of the aggregated Set Attention Block (SAB) weights culminates in the Multi-locus Attribution-Weighted Interaction (MAWI) Map.This approach quantifies the synergistic coordination between mutations, moving beyond single-variant associations to decompose the polygenic architecture of resistance. The interaction heatmap for Rifampicin (RIF) reveals a highly interconnected network with high signal density (Figure 17). This structure is characterized by "hub" variants, such as 3829770_T>>C and 1307598_C>>G, which exhibit strong cross-attention with multiple secondary mutations. This hub-driven topology quantitatively supports the hypothesis of epistatic fitness compensation, where robust resistance relies on a coordination of compensatory pathways to mitigate the fitness costs of primary mutations.
Figure 17: Interpretable attention heatmap of variant-variant interactions for Rifampicin (RIF) resistance. The matrix visualizes the self-attention weights derived from the Set Attention Block (SAB), representing the perceived strength of interaction between genomic variants. Darker red clusters identify "hub variants" that exert high influence over the collective resistance profile. These patterns reveal that VAMP-Net identifies non-linear epistatic relationships beyond simple additive effects, aligning with the biological understanding that complex resistance phenotypes often emerge from the co-occurrence of specific compensatory and primary mutations.
In contrast, the Rifabutin (RFB) network exhibits a sparse, localized architecture. The attention landscape is dominated by a few high-leverage interactions involving variants like 1070702_T>>C, set against a low-interaction background. This sparsity indicates that RFB resistance is governed by a select few non-additive synergistic effects that are mechanistically determinative, rather than the diffuse polygenic mechanism observed in RIF.
The contrasting patterns between these two rifamycins suggest that while they share target genes, their resistance modules are distinct. VAMP-Net’s ability to pinpoint these drug-specific interaction hubs provides an interpretable, auditable trace of the model’s reasoning. These results confirm known biological dependencies while highlighting novel epistatic interaction points that could serve as targets for future functional validation and the design of more effective combination therapies.
To confirm that the high-attention mutation pairs identified by the SAB transformer reflect genuine genomic dependencies rather than modeling artifacts, we performed a population-level validation using Fisher’s exact test and Odds Ratio (OR) analysis (Table 17). This approach provides an external statistical benchmark for the non-random dependencies captured by the attention mechanism. As detailed in Table Table 17, the statistical results for the primary interaction candidates demonstrate robust data-driven support:
  • RIF-associated pair ($1276232:C{>}A\leftrightarrow 3600576:C{>}T$): The combination of an infinite OR and an extremely low $p$-value ($9.296\times 10^{-89}$) indicates a highly structured co-occurrence. This suggests possible genetic linkage or strong co-selection under drug pressure, where the presence of one mutation is effectively determinative of the other.
  • RFB-associated pair ($1070702:T{>}C\leftrightarrow 985287:G{>}A$): The elevated odds ratio of $8.18$ ($p<0.01$) indicates significant enrichment, confirming that these mutations co-occur substantially more frequently than expected under independence.
These findings provide a critical layer of interpretability for VAMP-Net. By demonstrating that the highest-weighted interaction hubs are backed by significant population-level enrichment, we confirm that the model recovers mutation pairs consistent with the underlying structure of the M. tuberculosis population. This validates the attention heatmap as an auditable biological trace, effectively pinpointing non-random dependencies that may serve as targets for further functional investigation into epistatic fitness or shared resistance backgrounds.
Table 17: Statistical validation of representative high-attention mutation pairs using Fisher’s exact test. Only mutation pairs with strong statistical support are shown.
Drug Mutation pair OR $p$-value Interpretation
RIF $1276232:{C{>}A}\leftrightarrow 3600576:{C{>}T}$ $\infty$ $9.296\times 10^{-89}$ Extremely strong non-random co-occurrence; consistent with tight genetic linkage, co-selection, or a shared resistance background.
RFB $1070702:{T{>}C}\leftrightarrow 985287:{G{>}A}$ 8.18 $7.175\times 10^{-3}$ Significant enrichment, indicating a robust association beyond chance.
4.5.4 VCF Feature Channel Analysis
To validate the multi-channel design of the Quality-Aware 1D-CNN pathway, we quantified the contribution of eight VCF features through systematic ablation (Figure 18). The analysis confirms that the model transcends simple variant presence (GT) by leveraging granular contextual metrics:
@@ 4.5.4 VCF Feature Channel Analysis
  • Dominant Metric (FRS): The Fraction of Supporting Reads (FRS) is the most influential feature across both RIF and RFB models. This confirms the pathway effectively filters low-support variants, aligning with clinical standards for calling clonal mutations.
@@
  • Contextual Confidence: GT_CONF_PERCENTILE serves as the second most critical feature. By prioritizing relative confidence ranks over raw GT_CONF values, the model demonstrates a sophisticated capacity to utilize sample-specific comparative context.
  • Redundancy of Raw Metrics: The near-zero scores for Read_Depth (DP) and Genotype (GT) suggest that their information is successfully subsumed by normalized, derived features.
@@
These results demonstrate that the 1D-CNN pathway successfully extracts technically meaningful information, enabling a robust distinction between true resistance-conferring variants and sequencing noise.
Figure 18: Quantitative importance of technical quality metrics across EMB, INH, RIF, and RFB resistance prediction. The upper panel compares EMB (orange) and INH (blue), while the lower panel compares RIF (orange) and RFB (green). Across phenotypes, Fraction of Supporting Reads (FRS) and Genotype Confidence Percentile (GT_CONF_PERCENTILE) emerge as dominant technical predictors, indicating that VAMP-Net consistently leverages sequencing-confidence cues to recalibrate resistance predictions.
4.6 Discussion and Limitations
4.6.1 Architectural Efficacy and Integration Rationale
VAMP-Net demonstrates exceptional predictive fidelity across the anti-TB drug cohort, specifically achieving peak ROC-AUCs of 0.975 (RFB) and 0.970 (INH). The remarkably low variance across experimental seeds (std << 0.02) confirms that this performance is a robust consequence of the model’s structural design rather than stochastic initialization. Even in the presence of significant class imbalance, such as with Ethambutol (19.8% resistant), the model achieved the cohort’s highest PR-AUC (0.977), proving its resilience to the "curse of dimensionality" and skewed genomic datasets common in clinical settings. Comparative benchmarking reveals a significant performance gap between VAMP-Net and traditional baselines, with the MLP failing to exceed 85.88% accuracy. While the CNN effectively captured local spatial patterns, its higher false-negative rate (10.6%) relative to VAMP-Net (7.1%) underscores the necessity of attention mechanisms for modeling long-range epistatic interactions. VAMP-Net’s superior sensitivity in identifying 92.9% of resistant isolates while maintaining high specificity (91.3%) validates its utility as a high-precision diagnostic tool for complex antimicrobial resistance landscapes. Ultimately, the model’s success is rooted in its late-stage adaptive fusion philosophy. In contrast to Early Fusion, which suffered from a catastrophic specificity deficit (0.425) due to premature noise propagation, VAMP-Net’s multi-path design allows for modality-specific representational learning. By postponing integration until after contextual encoding, the framework achieved transformative gains, including a +48.8% increase in Specificity and a +30.9% increase in ROC-AUC. These results provide a mathematically grounded justification for VAMP-Net as a stable, noise-resilient framework for rapid molecular-based tuberculosis diagnostics.
4.6.2 VCF-Aware Recalibration and Clinical Utility
The granular case studies presented in Table 11 provide a critical window into the decision-making process of VAMP-Net, highlighting the Path-2 branch as a dynamic reliability filter. The substantial magnitude of observed probability shifts ($\Delta P\approx$ - 0.86) indicates that VCF-level metadata does not merely refine the genomic signal but fundamentally re-weights evidence based on technical fidelity. In Corrected cases, while Path-1 identifies resistance-like mutation patterns, Path-2 detects a conflict where weak sequencing support, characterized by low GT_CONF_PERCENTILE or insufficient read depth, renders the initial call unreliable. Conversely, in Amplified cases, Path-2 acts as a force-multiplier for high-quality data, shifting borderline predictions ($P(R)\approx$ 0.55) to high-confidence resistant calls (>> 0.90). By distinguishing between true biological polymorphisms and sequencing artifacts, VAMP-Net effectively treats variant presence as a necessary, but not sufficient, condition for resistance prediction. Beyond raw accuracy, VAMP-Net addresses the practical barriers to AI adoption in healthcare: calibration and operational efficiency. Brier score analysis confirms that VAMP-Net is a well-calibrated classifier, particularly for INH ($\sim 0.05$). High calibration ensures that predicted probabilities reflect the true likelihood of resistance, thereby reducing the clinical risk of suboptimal patient treatment. Furthermore, the architecture maintains a high-speed inference rate of $\sim 490$ samples per second. This computational efficiency, combined with high reproducibility and the reduction of grey-zone predictions through Path-2 recalibration, positions VAMP-Net as a reliable foundational tool for the next generation of rapid, molecular-based tuberculosis diagnostics.
4.6.3 Interpretability and Mechanistic Discovery
VAMP-Net’s dual-path architecture provides a mathematically grounded framework for identifying high-impact resistance mechanisms by effectively bridging high-dimensional genomics with biological and technical insights. Feature attribution analysis successfully recovered canonical resistance signatures, such as rpoB variants for RIF and RFB (e.g., 761140_A>>T; Score: 6.896), while internalizing complex co-selection patterns in MDR-TB strains by prioritizing cross-drug markers like katG (Score: 0.446). Beyond known targets, the Set Attention Block ($SAB$) identified novel, high-effect candidates—including the mmuM locus for INH (Score: 8.195) and the mmpL5/ppsD axis for RFB—which functional enrichment confirms constitute a statistically significant ($p=0.00239$) metabolic module centered on cell-wall remodeling and PDIM biosynthesis. Furthermore, systematic ablation of VCF features validates the 1D-CNN’s role as a signal-to-noise regulator that prioritizes the Fraction of Supporting Reads (FRS) and relative confidence rankings over raw depth. This "integrated audit" ensures that final predictions are governed by a synergy of genomic pathogenicity and technical reliability, allowing the model to dynamically down-weight biologically relevant variants that lack sufficient sequencing support. Collectively, these findings position VAMP-Net as an auditable discovery tool capable of reconstructing the multi-layered defense strategy of M. tuberculosis through the lens of both evolutionary selection and clinical data quality.
4.6.4 Limitations of the Study
While VAMP-Net demonstrates high predictive and interpretability performance, its scope is bounded by several factors. First, although our tokenization strategy effectively captures rare variants, generalizability remains constrained by the training cohort’s diversity; emerging mutations or rare lineages absent from the dataset may degrade performance. Second, while the 1D-CNN pathway successfully filters technical noise, it does not currently address broader systemic errors such as metagenomic contamination or sample-swapping in clinical workflows. Finally, the attention-based feature attribution provides a computational "causality hypothesis" rather than biochemical proof. The identified epistatic interactions and novel loci, though statistically significant, require experimental validation through site-directed mutagenesis or functional assays to confirm their mechanistic roles